can you please go step by step how to work these problems out but leaving the answer as a fraction!
4(3n+5)-2(2-4n)=6-2n

Answer by rothauserc(4718)

4(3n+5)-2(2-4n)=6-2n
12n+20 -4+8n = 6-2n
20n +16 = 6-2n
22n = -10
n = -10/22
n = -5/11